OpenStudy (anonymous):

5x+3y-5z=7 3x-3y+3z=3 6x-4y+6z=14

OpenStudy (anonymous):



OpenStudy (anonymous):

now im stuck!

OpenStudy (anonymous):

come on you got it; only two more to go

OpenStudy (anonymous):

i know but i tried everything.

OpenStudy (anonymous):

-12x+12y-12z=12 18x - 12y+18z=42 ------------------------ 6x +6z=54

OpenStudy (anonymous):

oh i found my mistake i think its suppose to be -12 =42=30

OpenStudy (anonymous):

(8x - 2z =10)*3 6x + 6z =30

OpenStudy (anonymous):

yeah i just did that to. i got x=2

OpenStudy (anonymous):

24x-6z=30 6x+6z=30 ------------ 30x=60 x=2

OpenStudy (anonymous):

z=3

OpenStudy (anonymous):

(2,4,3)

OpenStudy (anonymous):

i need to take a 30 minutes quiz its 5 problems can you help me?

OpenStudy (anonymous):

i can if you wait for 2 hrs

OpenStudy (anonymous):

when do you have to take it

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

wanna wait?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, see u in about 2 hrs